Use the information below and a 35% tax rate to calculate the WACC.

Common Stock Bonds
Units Outstanding 50,000 800
Market Price $14 $950
Cost 14% 7%

Please show written work and formula/s used, thank you :)
Answer #1

WACC = Weight of debt * Cost of debt * (1 - tax) + Weight of common equity * Cost of common equity

Value of debt = 800 * $950 = $760,000

Value of Equity = 50,000 * $14 = $700,000

Weight of debt = $760,000/($760,000 + $700,000) = 0.5205

Weight of equity = $700,000/($760,000 + $700,000) = 0.4795

WACC = 0.5205 * 7% * (1 - 35%) + 0.4795 * 14% = 9.08%